Guest User


a guest
Mar 16th, 2014
Not a member of Pastebin yet? Sign Up, it unlocks many cool features!
text 4.56 KB | None | 0 0
  1. Projects/SEM
  2. [edit] Scanning Electron Microscope
  4. We acquired a Cambridge Stereoscan 200 scanning electron microscope from a NASA contractor. The microscope arrived in not-running condition. Various repairs will need to take place before we have any hope of examining our first sample.
  5. Contents
  7. 1 Scanning Electron Microscope
  8. 1.1 Repair Log
  9. 1.1.1 August 5, 2012
  10. 1.1.2 August 8, 2012
  11. 1.1.3 August 11, 2012
  12. 1.1.4 ... thru 1 Sep 2012
  13. 1.2 Issues
  14. 1.3 Parts ordered
  16. [edit] Repair Log
  17. [edit] August 5, 2012
  19. We finally took our first steps with the SEM. After removing and opening up the power distribution board that connects directly to mains, we vacuumed out some rodent debris and noticed that shifting around the power distribution board caused relays to click and the "Operate" button to light intermittently. Apparently there are a couple broken pins within one of the Molex connectors that attach to the board and wiggling the cable around results in temporary operation. Ken ordered some replacement pins from Newark and we will probably re-crimp all of them. It turns out the busted looking toggle switch on the roughing pump is actually completely bypassed for direct control so we don't have to worry about replacing it. It also looks like there is also some water damage from having been housed in the previous owner's garage for years. Hopefully it won't cause too much trouble.
  21. We also obtained some more documentation from a person running the same model SEM in their living room but we haven't found anything relevant to the view screen yet. At what point is the CRT supposed to light up? We seem to not get any image if the system is either is either in On or On+Operate modes. The screen looks a little strange- as if there are bubbles or small beads inside the glass.
  23. We are very excited to continue working on the SEM once our replacement parts arrive!
  24. [edit] August 8, 2012
  26. New pins have been crimped replacing the two broken connector pins for connecting to the power distribution board. Now the power distribution does not shut off with a bump of the cable.
  28. We now see video output from the console using the composite CCTV BNC port on the back of the unit connected to a television. The built-in monitor is still not powering on.
  30. There is some clicking sound behind the CRT for the photograph projector which we need to investigate. Ken also found a burnt tantalum capacitor on the input for a voltage regulator on a circuit board next to the turbo pump which probably explains the burning smell experienced when connecting the that board to the main power supply. There are also a couple flickering LEDs on the main console board in a small row of solidly-illuminated LEDs which seems questionable.
  32. We still need to look at the waveforms from the various rectifiers in the power supply and the various regulators housed elsewhere to verify they are stable and not contributing to unexpected behavior.
  33. [edit] August 11, 2012
  35. Removed projector CRT and control board temporarily to assist in cable rerouting. Rectifier board was removed and cleaned and cables were rerouted through desk and reseated. Chris found a damaged electrolytic capacitor on the input for a voltage regulator on the regulator 1 board that may be, at the very least, the cause of two flickering indicator LEDs on the same board. Replacement capacitors and nylon standoffs have been ordered and known faulty capacitors will be replaced.
  36. [edit] ... thru 1 Sep 2012
  38. We believe the power supplies are now fixed. Several capacitors were replaced.
  40. Capacitor on vacuum control pcb was replaced, missing jumper was added there, now the button to select vac/vent works.
  42. Several problems with the computer have been fixed. Rodent pee on the board rusted some IC socket pins. We replaced some sockets and a 74LS74 we found was not working. Now the data entry keyboard works. The computer also uses a plugin card which has the potential to become misaligned. Believe that adjusting that caused the bottom row of the front panel keyboard to start working again.
  43. [edit] Issues
  45. Discovered 1 Sep 2012
  47. High voltage connection to display CRT came uncrimped.
  49. Resistors on scan processor overheat & flame when the deflection coils are connected.
  51. FOCUS/COARSE switch has connection/mechanical issues.
  53. Manual LEVEL pot spins freely with no stop.
  55. Refurbish rotary vacuum pump
  57. Refurbish turbo vacuum pump
  59. Dessicator sand? Maybe we can simply put it in the oven?
  60. [edit] Parts ordered
  62. 100x MOLEX - 08-50-0032 - CONTACT, FEMALE, 30-22AWG, CRIMP
Add Comment
Please, Sign In to add comment